# Emergence of a second law of thermodynamics in isolated quantum systems

### v1

@@ I Introduction
Irreversible evolution observed in nature is characterized by the second law of thermodynamics [1, 2, 3, 4], and is thus quantitatively captured by the increase in entropy over time of an isolated system. From the perspective of microscopic physics, however, all laws are time-reversal invariant and, thus, if a process is allowed by these laws, its reverse process must be, too. This means that entropy could, in principle, also decrease. This apparent conflict between the reversibility of physics on the microscopic level and the irreversibility of macroscopic physics, sometimes called Loschmidt’s paradox, has been debated since the earliest days of classical statistical mechanics [5, 6, 7]. As for the entropy, for classical systems, Liouville’s theorem implies that the Gibbs entropy formula is constant under Hamiltonian evolution [8], and also for quantum systems, the unitary evolution generated by the Schrödinger equation preserves the von Neumann entropy. In this Letter, we address this issue from the perspective of quantum systems, determining the sense in which entropy can increase over time in isolated quantum systems.
@@
With regards to quantum mechanics, despite substantial theoretical arguments [20, 21, 22, 23, 24, 25, 26] and experimental evidence [27, 28, 29, 27] for the equilibration of isolated quantum systems, no consensus exists about how isolated quantum systems equilibrate in the sense of something akin to the second law, as unitary evolution preserves the von Neumann entropy. By using the modern theory of equilibration of isolated quantum systems [21, 22] we show how the second law of thermodynamics is satisfied on average. In this framework, the system is examined relative to an arbitrary observable represented by a self-adjoint operator $O=\sum_{i=1}^{r}O_{i}\Pi_{i}^{O}$, where $r$ is the number of possible measurement outcomes. Assuming finite Hilbert space dimension $d$ (with $r<d$) and identifying the measurement outcomes with macrostates introduces, in a sense which we will quantify, the necessary lack of information, and ensuing emergent irreversibility.
@@
as in [30]. Note that instead of the projective valued measure (PVM) associated with the observable $O$ as we use it in the main text, one can define the probability distribution $\vec{p}$, with respect to a positive operator valued measure (POVM) $(E_{i})_{i=1}^{r}$ and all the following results follow analogously (see Appendix II.3).
@@
where the average $\langle\cdot\rangle_{\infty}$ is with respect to uniformly random times $0\leq t<\infty$, and the right-hand side (expressed in terms of big-$\mathcal{O}$ notation) vanishes in the limit where the number of microstates $d_{\mathrm{eff}}=\textnormal{Tr}[\omega^{2}]$ is much larger than that of macrostates $r.$
@@ III Results
Starting with the von Neumann entropy, one common and sensible-seeming attempt to formulate a second law for increasing entropy is to look at the entropy of time-averaged states $\langle\rho(t)\rangle_{T}$, where we have that $S_{\mathrm{vN}}[\langle\rho(t)\rangle_{T}]<S_{\mathrm{vN}}[\omega]$ if we compare to the infinite time averaged state $\omega$ (as we discuss in some more detail in Appendix I.1). While on a formal level, the entropy increases and approaches that of the infinite time average, such an argument fails to capture that in reality, an observer accesses the instantaneous state $\rho(t)$ at a well-defined time $t$, not the time averaged state $\langle\rho(t)\rangle_{T}$. It is this time parameter that progresses and with respect to which we observe equilibration, not the duration $T$ of the averaging window. For these reasons we examine how the Shannon observable entropy of the instantaneous state $\rho(t)$ behaves and whether this quantity satisfies the second law. By using Jensen’s inequality (see Lemma II.1 in Appendix B) we can arrive at a preliminary statement reminiscent of the second law:
@@
where $\varepsilon>0$ is an arbitrary energy parameter as in (7) (see Lemma II.1 in Appendix B for details). Inequality (10) improves previous bounds in Refs. [22, 23] quadratically in terms of their scaling with $r$, showing that in fact $r$ and $d_{\mathrm{eff}}$ appear with the same exponent.
@@
A statement akin to Theorem 1 can be derived for the observational entropy, which we discuss in further detail in Appendix I.2. For the observational entropy, it holds that
@@
The proof consists of two steps: (i) The first step relates a difference in Shannon entropies to a difference in state populations. The key here is that if the average distance between $\vec{p}_{\omega}$ and $\vec{p}(t)$ is small, then the average distance between their Shannon entropies $S_{\mathrm{Sh}}^{O}[\omega]$ and $S_{\mathrm{Sh}}^{O}\left[\rho(t)\right]$ should also be small. This connection can be made by combining the continuity bound for the Shannon entropy from [40] for the non-averaged quantities together with the result from Eq. (10). (ii) Secondly, we show that under the conditions described above, the population differences are small on average. This is the result from Eq. (10) and we prove it separately in Lemma II.1 in Appendix B. Both steps together prove the Theorem, and are described in detail in Appendix B. ∎
@@
$\delta(r,d_{\mathrm{eff}})=\frac{\log(1+r)}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}+H_{2}\left[\frac{1}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}\right],$ (14)
@@
This is a consequence of the generalized Chebyshev inequality $P[X\geq\kappa]\leq\frac{E[g(X)]}{g(\kappa)}$, where $X$ is a positive random variable and $g:\mathbb{R}\rightarrow\mathbb{R}$ a monotonously increasing, strictly positive function. Setting $X=\left|S_{\mathrm{sh}}[\rho(t)]-S_{\mathrm{sh}}[\omega]\right|$ and $\kappa:=\sqrt{\delta}$ completes the proof upon applying Eq. (11) from Theorem 1. For details, we refer the reader to Appendix II.2. ∎
The fluctuation bound in Corollary III has an interpretation in the sense of typicality [46, 47, 48]. In information theory, when sampling outputs from a fixed source, the weak Law of Large Numbers ensures that for large enough samples, the likely samples are those whose entropy is close to that of the source. This fluctuation bound can be interpreted from this perspective, offering a connection between information theory and thermodynamics. In this context, we are sampling at random from a fixed source – the expectation values $\vec{p}$ given by $\textnormal{Tr}[O\rho(t)]$ – and find that at all times it is likely that the entropy of this distribution of expectation values is close to the entropy of the distribution of expectation values of the equilibrium state. As such we may say that the equilibrium state $\omega$ is a ‘typical’ state when sampling expectation values of a state relative to an observable $O$ which equilibrates, and by extension states close to the equilibrium state in $S_{\mathrm{Sh}}^{O}[\,\cdot\,]$ are also typical.
@@
Though we discussed entropy defined relative to an observable which can be represented by a self-adjoint operator $O$, one can instead consider a more general positive operator valued measure (POVM) $M=(E_{i})_{i=1,\dots,r}$ and denote the corresponding probabilities $p_{i}(t)=\textnormal{Tr}[E_{i}\rho(t)]$. In Appendix II.3, we show that Theorem 1 and Corollary III readily generalize to this case without modification.
While this work focused on the Shannon observable entropy, we have proved analogous statements for the observational entropy in Appendix I.2, showing that the condition $r/d_{\mathrm{eff}}\ll 1$ is not only sufficient for equilibration of the Shannon observable entropy but also for the observational entropy. Still, one can construct examples where the Shannon observable entropy fluctuates on the order of $\log r$ and thus does not equilibrate, whereas the observational entropy’s fluctuations are relatively small because of the additional Boltzmann term which is of order $\log d$. This could lead one to conclude that the observational entropy equilibrates, while the Shannon observable entropy does not. Such a conclusion misses the point of having used an observable-relative notion of entropy in the first place – it is the observable whose equilibration is in question. If we are unable to distinguish fluctuations on the order of $\log r$ in the measurement outcome uncertainty, then we can simply choose a coarse-grained observable to begin with, where such fluctuations are operationally undetectable, and consider the Shannon entropy with respect to that observable. That being said, regardless of its role in understanding equilibration, the observational entropy does quantify the ignorance resulting from multiple states being compatible with a given measurement outcome [50], which the Shannon observable entropy alone does not capture.
Given that the ratio of observable outcomes to effective dimension in the case of a small subsystem of a much larger total system is generally favorable to equilibration (the former being bounded by the dimensionality of the subsystem), our results can likely be used to constrain the behavior of the von Neumann entropy of a subsystem. Since our focus was on the overall behavior of isolated systems, we have not addressed this here. Additionally, one may ask how our analysis generalizes to the context of other notions of entropy such as the family of Tsallis or Rényi entropies [51, 52, 53], by defining them relative to an observable in analogy to how we define the Shannon observable entropy. By combining entropy continuity bounds such as those of e.g. [52, 54] together with Lemma II.1 for PVMs or Lemma II.3 for POVMs (see Appendix), respective formulation of the second law might be obtained for such entropies.
@@ Acknowledgements.
The authors thank Pharnam Bakshinezhad, Felix Binder and Sandu Popescu for valuable discussions. T.R. thanks Sophie Engineer for assistance with programming. J.X. and M.H. would like to acknowledge funding from the European Research Council (Consolidator grant ‘Cocoquest’ 101043705). T.D. acknowledges support from ÖAW-JESH-Programme and the Brazilian agencies CNPq (Grant No. 441774/2023-7 and 200013/2024-6) as well as INCT-IQ through the project (465469/2014-0). This project is co-funded by the European Union (Quantum Flagship project ASPECTS, Grant Agreement No. 101080167). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union, REA or UKRI. Neither the European Union nor UKRI can be held responsible for them. This publication was made possible through the support of Grant 62423 from the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.
Supplemental material
Appendix A Additional entropies

### v2

@@ I Introduction
Irreversible evolution observed in nature is characterized by the second law of thermodynamics [1, 2, 3, 4], and is thus quantitatively captured by the increase in entropy over time of an isolated system. From the perspective of microscopic physics, however, all laws are time-reversal invariant and, thus, if a process is allowed by these laws, its reverse process must be, too. This means that entropy could, in principle, also decrease. This apparent conflict between the reversibility of physics on the microscopic level and the irreversibility of macroscopic physics, sometimes called Loschmidt’s paradox, has been debated since the earliest days of classical statistical mechanics [5, 6, 7, 8]. As for the entropy, for classical systems, Liouville’s theorem implies that the Gibbs entropy formula is constant under Hamiltonian evolution [9], and also for quantum systems, the unitary evolution generated by the Schrödinger equation preserves the von Neumann entropy. In this Letter, we address this issue from the perspective of quantum systems, determining the sense in which entropy can increase over time in isolated quantum systems.
@@
With regards to quantum mechanics, despite substantial theoretical arguments [21, 22, 23, 24, 25, 26, 27, 28] and experimental evidence [29, 30, 31, 29] for the equilibration of isolated quantum systems, no consensus exists about how isolated quantum systems equilibrate in the sense of something akin to the second law, as unitary evolution preserves the von Neumann entropy. By using the modern theory of equilibration of isolated quantum systems [22, 23] we show how the second law of thermodynamics is satisfied on average. In this framework, the system is examined relative to an arbitrary observable represented by a self-adjoint operator $O=\sum_{i=1}^{r}O_{i}\Pi_{i}^{O}$, where $r$ is the number of possible measurement outcomes. Assuming finite Hilbert space dimension $d$ (with $r<d$) and identifying the measurement outcomes with macrostates introduces, in a sense which we will quantify, the necessary lack of information, and ensuing emergent irreversibility.
@@
as in [32]. Note that instead of the projective valued measure (PVM) associated with the observable $O$ as we use it in the main text, one can define the probability distribution $\vec{p}$, with respect to a positive operator valued measure (POVM) $(E_{i})_{i=1}^{r}$ and all the following results follow analogously (see Appendix C).
@@
where the average $\langle\cdot\rangle_{\infty}$ is with respect to uniformly random times $0\leq t<\infty$, and the right-hand side (expressed in terms of big-$\mathcal{O}$ notation) vanishes in the limit where the number of microstates $d_{\mathrm{eff}}=1/\textnormal{Tr}[\omega^{2}]$ is much larger than that of macrostates $r.$
@@ III Results
Starting with the von Neumann entropy, one common and sensible-seeming attempt to formulate a second law for increasing entropy is to look at the entropy of time-averaged states $\langle\rho(t)\rangle_{T}$, where we have that $S_{\mathrm{vN}}[\langle\rho(t)\rangle_{T}]<S_{\mathrm{vN}}[\omega]$ if we compare to the infinite time averaged state $\omega$ (as we discuss in some more detail in Appendix II.1). While on a formal level, the entropy increases and approaches that of the infinite time average, such an argument fails to capture that in reality, an observer accesses the instantaneous state $\rho(t)$ at a well-defined time $t$, not the time averaged state $\langle\rho(t)\rangle_{T}$. It is this time parameter that progresses and with respect to which we observe equilibration, not the duration $T$ of the averaging window. For these reasons we examine how the Shannon observable entropy of the instantaneous state $\rho(t)$ behaves and whether this quantity satisfies the second law. By using Jensen’s inequality (see Lemma I.1 in Appendix A) we can arrive at a preliminary statement reminiscent of the second law:
@@
where $\varepsilon>0$ is an arbitrary energy parameter as in (7) (see Lemma I.1 in Appendix A for details). Inequality (10) improves previous bounds in Refs. [23, 24] quadratically in terms of their scaling with $r$, showing that in fact $r$ and $d_{\mathrm{eff}}$ appear with the same exponent.
@@
$\left\langle\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\right\rangle_{T}\leq\log(r-1)\eta_{\varepsilon,T}+H_{2}[\eta_{\varepsilon,T}],$ (11)
@@
A statement akin to Theorem 2 can be derived for the observational entropy, which we discuss in further detail in Appendix II.2. For the observational entropy, it holds that
@@
The proof consists of two steps: (i) The first step relates a difference in Shannon entropies to a difference in state populations. The key here is that if the average distance between $\vec{p}_{\omega}$ and $\vec{p}(t)$ is small, then the average distance between their Shannon entropies $S_{\mathrm{Sh}}^{O}[\omega]$ and $S_{\mathrm{Sh}}^{O}\left[\rho(t)\right]$ should also be small. This connection can be made by combining the continuity bound for the Shannon entropy from [42] for the non-averaged quantities together with the result from Eq. (10). (ii) Secondly, we show that under the conditions described above, the population differences are small on average. This is the result from Eq. (10) and we prove it separately in Lemma I.1 in Appendix A. Both steps together prove the Theorem, and are described in detail in Appendix A. ∎
@@
$\delta(r,d_{\mathrm{eff}})=\frac{\log(r-1)}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}+H_{2}\left[\frac{1}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}\right],$ (14)
@@
This is a consequence of the generalized Chebyshev inequality $P[X\geq\kappa]\leq{E[g(X)]}/{g(\kappa)}$, where $X$ is a positive random variable and $g:\mathbb{R}\rightarrow\mathbb{R}$ a monotonously increasing, strictly positive function. Setting $X=\left|S_{\mathrm{sh}}[\rho(t)]-S_{\mathrm{sh}}[\omega]\right|$ and $\kappa:=\sqrt{\delta}$ completes the proof upon applying Eq. (26) from Theorem 2. For details, we refer the reader to Appendix I.2. ∎
The fluctuation bound in Corollary III has an interpretation in the sense of typicality [48, 49, 50, 27]. In information theory, when sampling outputs from a fixed source, the weak Law of Large Numbers ensures that for large enough samples, the likely samples are those whose entropy is close to that of the source. This fluctuation bound can be interpreted from this perspective, offering a connection between information theory and thermodynamics. In this context, we are sampling at random from a fixed source – the expectation values $\vec{p}$ given by $\textnormal{Tr}[O\rho(t)]$ – and find that at all times it is likely that the entropy of this distribution of expectation values is close to the entropy of the distribution of expectation values of the equilibrium state. As such we may say that the equilibrium state $\omega$ is a ‘typical’ state when sampling expectation values of a state relative to an observable $O$ which equilibrates, and by extension states close to the equilibrium state in $S_{\mathrm{Sh}}^{O}[\,\cdot\,]$ are also typical.
@@ IV Discussion
Though we discussed entropy defined relative to an observable which can be represented by a self-adjoint operator $O$, one can instead consider a more general positive operator valued measure (POVM) $M=(E_{i})_{i=1,\dots,r}$ and denote the corresponding probabilities $p_{i}(t)=\textnormal{Tr}[E_{i}\rho(t)]$. In Appendix C, we show that Theorem 2 and Corollary III readily generalize to this case without modification.
While this work focused on the Shannon observable entropy, we have proved analogous statements for the observational entropy in Appendix II.2, showing that the condition $r/d_{\mathrm{eff}}\ll 1$ is not only sufficient for equilibration of the Shannon observable entropy but also for the observational entropy. Still, one can construct examples where the Shannon observable entropy fluctuates on the order of $\log r$ and thus does not equilibrate, whereas the observational entropy’s fluctuations are relatively small because of the additional Boltzmann term which is of order $\log d$. This could lead one to conclude that the observable equilibrates while in fact it does not, as the Shannon entropy shows. Regardless of its role in understanding equilibration, the observational entropy does quantify the ignorance resulting from multiple states being compatible with a given measurement outcome [52], which the Shannon observable entropy alone does not capture. As we have shown, this many-to-one relationship between states and outcomes instead dictates the dynamics of the Shannon entropy, depending on the dynamical participation of those states as encoded in $d_{\mathrm{eff}}$.
Given that the ratio of observable outcomes to effective dimension in the case of a small subsystem of a much larger total system is generally favorable to equilibration (the former being bounded by the dimensionality of the subsystem), our results can likely be used to constrain the behavior of the von Neumann entropy of a subsystem. Since our focus was on the overall behavior of isolated systems, we have not addressed this here. Additionally, one may ask how our analysis generalizes to the context of other notions of entropy such as the family of Tsallis or Rényi entropies [53, 54, 55], by defining them relative to an observable in analogy to how we define the Shannon observable entropy. By combining entropy continuity bounds such as those of e.g. [54, 56] together with Lemma I.1 for PVMs or Lemma C for POVMs (see Appendix), respective formulation of the second law might be obtained for such entropies.
@@ Acknowledgements.
[…]
Appendices
Appendix A Proofs for entropy bounds
In this section, we prove all our main results and provide additional background to the statements and the formalism we use. In Sec. I.1 we provide the proofs for the average entropy difference bounds, that is Theorem 2 and Proposition 3. Following this, Sec. I.2 contains the proofs for the fluctuation bounds of Corollary III and Corollary 3.
I.1 Proofs for averaged entropy differences
Before discussion of the proof of Theorem 2, more details regarding the inequality (9) are supplied, in particular regarding the applicability of Jensen’s inequality: {lemma}[] Given $\rho(t)$, $\omega$ and $O$ as defined in the main text, we have in the infinite time-average,
Proof.
The proof of this statement is an application of the Jensen inequality which states that for a concave, real-valued function $g[X]$ the following inequality holds: $\left\langle g[X]\right\rangle\leq g\left[\left\langle X\right\rangle\right]$. In this context, $X$ is a real-valued random variable that can in general be defined on a convex subset of $\mathbb{R}^{r}$, as shown for example in Theorem 7.11 of [61]. For our setting, we have that $g[\,\cdot\,]=S_{\mathrm{Sh}}^{O}[\,\cdot\,]$ is concave as a function of the probability vector $\vec{p}(t)$ defined by the observable outcomes. Note that it is not concave as a function of $t$, though this is, of course, not needed here. Formally, the Shannon observable entropy as defined in the main text is a function of density matrices, but since it only depends on the population vectors, we simplify the notation here to simply write $S_{\mathrm{Sh}}^{O}[\vec{p}]$ to actually mean $S_{\mathrm{Sh}}^{O}[\rho]$, where $\vec{p}$ is the population vector of $\rho$ relative to the observable $O$. By identifying $X=\vec{p}(t)$, and taking the uniform measure over the space $t\in[0,T]$, Jensen’s inequality reads
[…]
Moving on towards proving our main result, Theorem 2, we devise the following proof strategy:
We start with the first part. Let us recall that $\vec{p}(t)=(p_{1},\dots,p_{r})$ with $p_{i}=\textnormal{Tr}[\Pi_{i}^{O}\rho]$ is the vector of populations of $\rho(t)$ with respect to the observable’s eigenspaces. Similarly, $\vec{p}_{\omega}$ is defined for the infinite time averaged state $\omega$. Then, in Eq. (10) of the main text we claimed that $\langle\|\vec{p}(t)-\vec{p}_{\omega}\|_{1}\rangle_{T}$ is small, which we prove in detail as part of Lemma I.1. The key reason why the populations’ differences are small on average is based on the results for observables from the theory of equilibration on average, which state that $\textnormal{Tr}[O\rho(t)]$ is close to $\textnormal{Tr}[O\omega]$ most of the time. One subtle point here concerns the number of macrostates: while for the observables’ expectation values it was sufficient to have that $d_{\mathrm{eff}}$ is large for equilibration, now, the effective dimension has to be large compared to $r$, the number of macrostates corresponding to the observable $O$. We can therefore state that observable expectation values equilibrate under weaker conditions than the population vectors.
{lemma}
Proof.
$\left\langle\left|\textnormal{Tr}\left[\Pi_{i}^{O}\rho(t)\right]-\textnormal{Tr}\left[\Pi_{i}^{O}\omega\right]\right|^{2}\right\rangle_{T}$
$\quad\leq f(\varepsilon,T)\sqrt{\textnormal{Tr}\left[\Pi_{i}^{O\dagger}\Pi_{i}^{O}\omega^{2}\right]\textnormal{Tr}\left[\Pi_{i}^{O}\Pi_{i}^{O\dagger}\omega^{2}\right]}.$ (20)
[…]
$=f(\varepsilon,T)\textnormal{Tr}\left[\omega^{2}\right]$ (24)
$=\frac{f(\varepsilon,T)}{d_{\mathrm{eff}}}.$ (25)
Now that we have checked the first item of our proof strategy outlined in the beginning of Sec. A, we are ready to show Theorem 2, which we restate in the following for completeness.
Theorem 2 (Second law on average).
Proof.
We subdivide the proof of this statement into two steps: (1) We express the left-hand side of Eq. (26) in terms of population differences, giving us a bound on the entropy difference on average expressed as a function of the average difference in populations between $\rho(t)$ and the infinite-time averaged state $\omega$. (2) Then, we use the result from Lemma I.1, which we can then insert into the bound found in the first step, finalizing the proof of this Proposition.
Step (1).
$\|\vec{\vartheta}\|_{1}=\frac{1}{2}\sum_{i=1}^{r}|\vartheta_{i}|,$ (28)
$\left\langle\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\right\rangle_{T}$
$\qquad=\left\langle\left|S_{\mathrm{Sh}}[\vec{p}(t)]-S_{\mathrm{Sh}}[\vec{p}_{\omega}]\right|\right\rangle_{T}$ (29)
$\qquad\leq\log(r-1)\langle\|\vec{\vartheta}\|_{1}\rangle_{T}+\langle H_{2}[\|\vec{\vartheta}\|_{1}]\rangle_{T}.$ (30)
$\left\langle\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\right\rangle_{T}$
$\qquad\leq\log(r-1)\langle\|\vec{\vartheta}\|_{1}\rangle_{T}+H_{2}[\langle\|\vec{\vartheta}\|_{1}\rangle_{T}].$ (32)
Step (2).
Recalling that the binary entropy $H_{2}[x]$ is monotonously growing for $x\in[0,1/2]$ allows us to insert the inequality from Lemma I.1, $\langle\|\vec{\vartheta}\|_{1}\rangle_{T}\leq\eta_{\varepsilon,T}$ into the result from the first step (32). We find,
I.2 Proofs for fluctuation theorems
Here we provide a proof for the fluctuation bounds for the Shannon observable entropy (Corollary III), which can also by straightforward analogy serve as the proof of Corollary 3 concerning the fluctuations of the observational entropy.
See III
Proof.
We provide additional details for the proof of Corollary III by showing a simple proof for the generalized Chebyshev inequality, which we recall here for completeness:
$P[X\geq\kappa]\leq\frac{E[g(X)]}{g(\kappa)}.$ (34)
[…]
$E[X]=\frac{\log(1+r)}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}+H_{2}\left[\frac{1}{2}\sqrt{\frac{r}{d_{\mathrm{eff}}}}\right]=:\delta,$ (38)
as a direct consequence of Proposition 2. Setting $\kappa:=\sqrt{\delta}$ yields
which completes the proof as claimed in the main text. ∎
Instead of using the symmetric fluctuation bounds in Corollary III where both the magnitude of the fluctuations as well as the probability of them is compared to $\sqrt{\delta}$, one can use a more general re-scaled form. Starting with the Chebyshev inequality (35) we could instead write a more general bound
$P\left[\left|S_{\mathrm{Sh}}^{O}[\rho(t)]-S_{\mathrm{Sh}}^{O}[\omega]\right|\geq\mu\right]\leq\frac{\delta}{\mu},$ (40)
where on the left-hand side we consider fluctuations of the Shannon entropy of the order of $\mu$. At the same time, on the right-hand side, the inverse is multiplied giving $\delta/\mu$. This generalized bound may be useful in the setting where one considers entropy fluctuations of a specific magnitude $\mu>\sqrt{\delta}$ and therefore the bound in Corollary III does not directly apply anymore. The Chebyshev inequality is in general not tight, though, and therefore it is expected that for some choices of parameters, the probability of fluctuations is overestimated by this method.
Appendix B Additional entropies